OpenStudy (anonymous):

please help......the x-intercept of a line is (3/2,0) and the line slope =4, what is the equation of the line in slope-intercept form?

OpenStudy (aum):

y = mx + b is the general slope intercept form. Slope m = 4 y = 4x + b Point (3/2, 0) is on the line. Put x = 3/2, y = 0 and solve for b.

OpenStudy (anonymous):

so 0=4(3/2)+b?

OpenStudy (anonymous):

i got -6

OpenStudy (aum):

Yes. Put b back in y = 4x + b and you will have your answer.

OpenStudy (anonymous):

so ill put -6?

OpenStudy (anonymous):

where b is?

OpenStudy (aum):

Put b = -6 in y = 4x + b.

OpenStudy (anonymous):

and that would be the answer

OpenStudy (aum):

Yes. y = 4x - 6 is the equation of the line in slope-intercept form.

OpenStudy (anonymous):

thank you so much

OpenStudy (aum):

You are welcome.